I have to write 2 equations and solve them. I do not understand how to write an equation
Answer by doukungfoo(195)

Well according to The American HeritageŽ Science Dictionary, An equation is a statement asserting the equality of two expressions, usually written as a linear array of symbols that are separated into left and right sides and joined by an equal sign.
An example of an equation is:
5 + 4 = 9
In algebra more often than not we have equations with variables and we are asked to solve the equation. This means to figure out what value or values the variable could be to make the equation true.
Example of equation with variable:
5 + x = 9
x = 9 - 5
x = 4
Proof:
5 + 4 = 9
9 = 9
The examples here are intended to show the basics of what an equation is. Equations can be very complex and can contain more than one variable. But the one thing that seperates an equation for other types of algebraic expressions is the equal sign. For an expression to be called an equation it must contain and equal sign.
hope this helps